# GSL-LPA: Fast Label Propagation Algorithm (LPA) for Community Detection with no Internally-Disconnected Communities

### v2

@@ Abstract.
Community detection is the problem of identifying tightly connected clusters of nodes within a network. Efficient parallel algorithms for this play a crucial role in various applications, especially as datasets expand to significant sizes. The Label Propagation Algorithm (LPA) is commonly employed for this purpose due to its ease of parallelization, rapid execution, and scalability. However, it may yield internally disconnected communities. This technical report introduces GSL-LPA, derived from our parallelization of LPA, namely GVE-LPA. Our experiments on a system with two 16-core Intel Xeon Gold 6226R processors show that GSL-LPA not only mitigates this issue but also surpasses FLPA, igraph LPA, and NetworKit LPA by $\times5$, 10,300×10,300$\times$10 , 300 ×, and 5.8×5.8$\times$5.8 ×, respectively, achieving a processing rate of $844M$ edges/s on a $3.8B$ edge graph. Additionally, GSL-LPA scales at a rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
@@ 1. Introduction
Community detection refers to the task of identifying groups of vertices characterized by dense internal connections and sparse connections between groups (Fortunato, 2010). These groups, often referred to as communities or clusters (Abbe, 2018), offer valuable insights into the structure and function of networks (Fortunato, 2010). The problem of community detection is NP-hard and finds applications in various domains, including topic discovery, protein annotation, recommendation systems, and targeted advertising (Gregory, 2010). An additional challenge arises from the lack of prior knowledge about the number and size distribution of communities (Blondel et al., 2008). Heuristic-based approaches are commonly employed for community detection (Clauset et al., 2004; Duch and Arenas, 2005; Reichardt and Bornholdt, 2006; Raghavan et al., 2007; Rosvall and Bergstrom, 2008; Blondel et al., 2008; Xie et al., 2011; Whang et al., 2013; Kloster and Gleich, 2014; Traag et al., 2019; You et al., 2020; Traag and Šubelj, 2023). The modularity metric (Newman, 2006a) is used to measure the quality of communities identified. The communities identified are considered intrinsic when based solely on network topology and disjoint when each vertex belongs to only one community (Gregory, 2010; Coscia et al., 2011).
@@
The Label Propagation Algorithm (LPA), also known as RAK (Raghavan et al., 2007), is a widely adopted diffusion-based heuristic for community detection, capable of identifying communities of moderate quality. It offers simplicity, speed, and scalability advantages over the Louvain method (Blondel et al., 2008), another effective community detection algorithm renowned for its high-quality results. Specifically, we observe LPA to outpace Louvain by 2.314×2.3-14$\times$2.3 - 14 × in terms of speed, while identifying communities of $3-30\%$ lower quality. LPA’s efficiency is attributed to its avoidance of repeated optimization steps and its ease of parallelization. Thus, LPA is well-suited for applications prioritizing high performance at the expense of slightly lower result quality. In our experimentation with alternative methods such as COPRA (Gregory, 2010), SLPA (Xie et al., 2011), and LabelRank (Xie et al., 2013), we found LPA to be the most efficient, while identifying communities of equivalent quality (Sahu, 2023d).
@@
However, LPA is susceptible to identifying internally disconnected communities. Through our experimental analysis, we observe that up to $3.1\%$, $0.3\%$, and $14.5\%$ of communities identified using FLPA (Traag and Šubelj, 2023), igraph LPA (Csardi et al., 2006), and NetworKit LPA (Staudt et al., 2016) exhibit this issue. Rectifying such disconnected communities is crucial to ensuring the accuracy and robustness of community detection algorithms. Several studies have addressed internally-disconnected communities as a post-processing step (Raghavan et al., 2007; Gregory, 2010; Hafez et al., 2014; Luecken, 2016; Wolf et al., 2019). This is also the methodology we adopt here.
1.1. Our Contributions
This report introduces GSL-LPA [^1] https://github.com/puzzlef/rak-communities-openmp, a parallel implementation of the Label Propagation Algorithm (LPA) designed to identify communities without internal disconnections. GSL-LPA is derived from our parallelization of LPA, referred to as GVE-LPA (Sahu, 2023c). Experimental evaluations conducted on a system equipped with two 16-core Intel Xeon Gold 6226R processors demonstrate that GSL-LPA not only resolves the issue of disconnected communities but also surpasses FLPA (Traag and Šubelj, 2023), igraph LPA (Csardi et al., 2006), and NetworKit LPA (Staudt et al., 2016) by $\times5$, 10,300×10,300$\times$10 , 300 ×, and 5.8×5.8$\times$5.8 ×, respectively. On a graph with $3.8B$ edges, GSL-LPA achieves a processing rate of $844M$ edges/s. Additionally, GSL-LPA scales at a rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
@@ 2. Related work
Communities within networks serve as functional units or meta-nodes (Chatterjee and Saha, 2019). Identifying these divisions in an unsupervised manner is crucial in various domains, including drug discovery, disease prediction, protein annotation, topic discovery, link prediction, recommendation systems, customer segmentation, inferring land use, and criminal identification (Karataş and Şahin, 2018). Numerous schemes have been devised for community detection (Clauset et al., 2004; Duch and Arenas, 2005; Reichardt and Bornholdt, 2006; Raghavan et al., 2007; Rosvall and Bergstrom, 2008; Blondel et al., 2008; Xie et al., 2011; Whang et al., 2013; Kloster and Gleich, 2014; Traag et al., 2019; You et al., 2020; Traag and Šubelj, 2023). These can be categorized into three basic approaches: bottom-up, top-down, and data-structure based (Souravlas et al., 2021). Additionally, they can be classified as divisive, agglomerative, or multi-level methods (Zarayeneh and Kalyanaraman, 2021). To evaluate these methods, fitness scores like modularity (Newman, 2006b) are used. Modularity measures the relative density of links within communities compared to those outside and ranges from $-0.5$ to $1$.
The Label Propagation Algorithm (LPA) is a diffusion-based method for identifying communities. Techniques for parallelizing LPA include vertex assignment with guided scheduling (Staudt and Meyerhenke, 2015), parallel bitonic sort (Soman and Narang, 2011), and pre-partitioning of the graph (Kuzmin et al., 2015).
In terms of quality, LPA may yield communities with low modularity score, resulting in the emergence of a dominant community that envelops the majority of the nodes, obscuring finer community structures (Gregory, 2010). To address instability of result, community structure quality, and performance due to vertex processing order, a few ordering strategies have been attempted. These include updating only the active nodes (Xie and Szymanski, 2011) or prioritizing specific subsets based on properties, like core or boundary nodes (Gui et al., 2018). In some cases, oscillations in label assignments may arise instead of convergence. To address this, asynchronous mode (Leung et al., 2009), alternating updates of independent node subsets (Cordasco and Gargano, 2012), and parallel graph coloring techniques (Cordasco and Gargano, 2012) have been attempted. Raghavan et al. (Raghavan et al., 2007) note that after just five iterations, labels of $95\%$ of nodes converge to their final values.
@@
A few open-source implementations and software packages have been developed for community detection utilizing the LPA. Fast Label Propagation Algorithm (FLPA) (Traag and Šubelj, 2023) is a variant of the LPA that employs a queue-based method, processing only vertices with recently updated neighborhoods. NetworKit (Staudt et al., 2016) is a software package built to analyze the structural attributes of graph datasets containing billions of connections. It integrates C++ kernels with a Python frontend, and features a parallel implementation of LPA. igraph (Csardi et al., 2006) is a similar package that is written in C, and offers Python, R, and Mathematica frontends. It enjoys widespread adoption in academic research and includes an LPA implementation.
Raghavan et al. (Raghavan et al., 2007) note that their Label Propagation Algorithm (LPA) for community detection can uncover internally-disconnected communities (Gregory, 2010). Genetic algorithms (Hesamipour et al., 2022) and expectation minimization/maximization algorithms (Ball et al., 2011; Hafez et al., 2014) can also detect such disconnected communities. To address this, they propose applying Breadth-First Search (BFS) to subnetworks in each individual group to separate the disconnected communities. This can be achieved with a time complexity of $O(M+N)$. Gregory (Gregory, 2010) extends LPA with the Community Overlap PRopagation Algorithm (COPRA), which removes nested communities and splits disconnected communities using a method similar to Raghavan et al. (Raghavan et al., 2007). Hafez et al. (Hafez et al., 2014) introduce a community detection algorithm using Bayesian Network and Expectation Minimization (BNEM), which includes a step to detect disconnected components within communities. If detected, disconnected components are assigned new community labels. This scenario arises when the network has more communities than the specified number $k$ in the algorithm. Hesamipour et al. (Hesamipour et al., 2022) propose a genetic algorithm for community detection, integrating similarity-based and modularity-based approaches. They use an MST-based representation to address issues like disconnected communities and ineffective mutations. Goel et al. (Goel et al., 2023) present a tangential work, where they identify a set of nodes in a network, known as a Structural Hole Spanner (SHS), that act as a bridge among different otherwise disconnected communities.
@@ 5.2. Comparing Performance of GSL-LPA
GSL-LPA demonstrates average speedups of $\times5$, 10,300×10,300$\times$10 , 300 ×, and 5.8×5.8$\times$5.8 × over FLPA, igraph LPA, and NetworKit LPA respectively. Specifically, on the sk-2005 graph, GSL-LPA identifies communities within $4.5$ seconds, achieving a processing rate of $844$ million edges/s. Furthermore, Figure 4(c) presents the modularity of communities obtained by each implementation. On average, GSL-LPA achieves modularity that is $7.1\%$ and $0.7\%$ higher than FLPA (particularly on road networks and protein k-mer graphs) and igraph LPA respectively. Additionally, it records a $3.6\%$ lower modularity compared to NetworKit LPA (especially on protein k-mer graphs).

### v3

@@ Abstract.
Community detection is the problem of identifying tightly connected clusters of nodes within a network. Efficient parallel algorithms for this play a crucial role in various applications, especially as datasets expand to significant sizes. The Label Propagation Algorithm (LPA) is commonly employed for this purpose due to its ease of parallelization, rapid execution, and scalability — however, it may yield internally disconnected communities. This technical report introduces GSL-LPA, derived from our parallelization of LPA, namely GVE-LPA. Our experiments on a system with two 16-core Intel Xeon Gold 6226R processors show that GSL-LPA not only mitigates this issue but also surpasses FLPA, igraph LPA, and NetworKit LPA by $\times5$, 10,300×10,300$\times$10 , 300 ×, and 5.8×5.8$\times$5.8 ×, respectively, achieving a processing rate of $844M$ edges/s on a $3.8B$ edge graph. Additionally, GSL-LPA scales at a rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
@@ 1. Introduction
Community detection refers to the task of identifying groups of vertices characterized by dense internal connections and sparse connections between groups (Fortunato, 2010). These groups, often referred to as communities or clusters (Abbe, 2018), offer valuable insights into the structure and function of networks (Fortunato, 2010). Community detection has a wide range of applications across multiple fields. In ecology, it is applied to assess if food webs are structured into compartments, where species within each compartment interact more frequently with each other than with those in different compartments (Guimerà et al., 2010). Within healthcare, it helps in analyzing the dynamics of populations vulnerable to epidemic diseases (Salathé and Jones, 2010), detecting conditions such as lung cancer (Bechtel et al., 2005), and classifying tumor types using genomic data (Haq and Wang, 2016). In drug discovery, community detection aids in identifying groups of similar compounds or target proteins, which supports the discovery of novel therapeutic agents (Ma et al., 2019). Furthermore, community detection has been effectively used to reveal the structure and evolution of complex biological networks, including metabolic networks (Kim et al., 2009), Gene Regulatory Networks (GRNs) (Rivera et al., 2010), and Lateral Gene Transfer (LGT) networks (Popa et al., 2011). It also plays a significant role in understanding human brain networks, enabling insights into their structural and functional organization (He and Evans, 2010).
@@
The problem of community detection is due to the lack of prior knowledge about the number of communities and their size distribution (Gregory, 2010). Heuristic-based approaches are commonly employed for community detection (Raghavan et al., 2007; Blondel et al., 2008; Xie et al., 2011; Traag and Šubelj, 2023; Clauset et al., 2004; Duch and Arenas, 2005; Reichardt and Bornholdt, 2006; Kloster and Gleich, 2014; Traag et al., 2019; You et al., 2020; Rosvall and Bergstrom, 2008; Whang et al., 2013). The modularity metric (Newman, 2006a) is used to measure the quality of communities identified. The communities identified are considered intrinsic when based solely on network topology and disjoint when each vertex belongs to only one community (Gregory, 2010; Coscia et al., 2011). The Label Propagation Algorithm (LPA), also known as RAK (Raghavan et al., 2007), is a widely adopted diffusion-based heuristic for community detection, capable of identifying communities of moderate quality. It offers simplicity, speed, and scalability advantages over the Louvain method (Blondel et al., 2008), another effective community detection algorithm renowned for its high-quality results. Specifically, we observe LPA to outpace Louvain by 2.314×2.3-14$\times$2.3 - 14 × in terms of speed, while identifying communities of $3-30\%$ lower quality. LPA’s efficiency is attributed to its avoidance of repeated optimization steps and its ease of parallelization. Thus, LPA is well-suited for applications prioritizing high performance at the expense of slightly lower result quality. While LPA typically yields communities with lower modularity scores, it has been observed to achieve high Normalized Mutual Information (NMI) score relative to the ground truth (Peng et al., 2014). In our experimentation with alternative label-propagation based methods such as COPRA (Gregory, 2010), SLPA (Xie et al., 2011), and LabelRank (Xie et al., 2013), we found LPA to be the most efficient, while identifying communities of equivalent quality (Sahu, 2023d).
@@
However, LPA is susceptible to identifying a large number of internally disconnected communities. Through our experimental analysis, we observe that up to $3.1\%$, $0.3\%$, and $14.5\%$ of communities identified using FLPA (Traag and Šubelj, 2023), igraph LPA (Csardi et al., 2006), and NetworKit LPA (Staudt et al., 2016) exhibit this issue. Rectifying such disconnected communities is crucial to ensuring the accuracy and robustness of community detection algorithms. While a number of studies have proposed addressing internally-disconnected communities as a post-processing step, they do not present a concrete algorithm to do so (Raghavan et al., 2007; Gregory, 2010; Hafez et al., 2014; Luecken, 2016; Wolf et al., 2019). In this technical report, we present a GSL-LPA, which is based upon an efficient multicore implementation of LPA (Sahu, 2023c), and utilizes our BFS-based approach for splitting internally-disconnected communities — which we show, outperforms other approaches for splitting the identified disconnected communities.
@@ 2. Related work
A number of schemes have been devised for community detection (Raghavan et al., 2007; Blondel et al., 2008; Xie et al., 2011; Traag and Šubelj, 2023; Clauset et al., 2004; Duch and Arenas, 2005; Reichardt and Bornholdt, 2006; Rosvall and Bergstrom, 2008; Whang et al., 2013; Kloster and Gleich, 2014; Traag et al., 2019; You et al., 2020). These can be categorized into three basic approaches: bottom-up, top-down, and data-structure based (Souravlas et al., 2021). Additionally, they can be classified as divisive, agglomerative, or multi-level methods (Zarayeneh and Kalyanaraman, 2021). To evaluate these methods, fitness scores like modularity (Newman, 2006b) are used.
The Label Propagation Algorithm (LPA) is a diffusion-based method for identifying communities. Techniques for parallelizing LPA include vertex assignment with guided scheduling (Staudt and Meyerhenke, 2015), parallel bitonic sort (Soman and Narang, 2011), and pre-partitioning of the graph (Kuzmin et al., 2015). In terms of quality, LPA may yield communities with low modularity score, resulting in the emergence of a dominant community that envelops the majority of the nodes, obscuring finer community structures (Gregory, 2010). To address instability of result, community structure quality, and performance due to vertex processing order, a few ordering strategies have been attempted. These include updating only the active nodes (Xie and Szymanski, 2011) or prioritizing specific subsets based on properties, like core or boundary nodes (Gui et al., 2018). In some cases, oscillations in label assignments may arise instead of convergence. To address this, asynchronous mode (Leung et al., 2009), alternating updates of independent node subsets (Cordasco and Gargano, 2012), and parallel graph coloring techniques (Cordasco and Gargano, 2012) have been attempted. Raghavan et al. (Raghavan et al., 2007) note that after just five iterations, labels of $95\%$ of nodes converge to their final values.
@@
A few open-source implementations and software packages have been developed for community detection utilizing the LPA. Fast Label Propagation Algorithm (FLPA) (Traag and Šubelj, 2023) is a variant of the LPA that employs a queue-based method, processing only vertices with recently updated neighborhoods. NetworKit (Staudt et al., 2016) is a software package built to analyze the structural attributes of graph datasets containing billions of connections. It integrates C++ kernels with a Python frontend, and features a parallel implementation of LPA. igraph (Csardi et al., 2006) is a similar package that is written in C, and offers Python, R, and Mathematica frontends. It has good adoption, and includes an LPA implementation.
Raghavan et al. (Raghavan et al., 2007) note that their Label Propagation Algorithm (LPA) for community detection can uncover internally-disconnected communities (Gregory, 2010). Genetic algorithms (Hesamipour et al., 2022) and expectation minimization/maximization algorithms (Ball et al., 2011; Hafez et al., 2014) can also detect such disconnected communities. To address this, they propose applying Breadth-First Search (BFS) to subnetworks in each individual group to separate the disconnected communities. This can be achieved with a time complexity of $O(M+N)$. Gregory (Gregory, 2010) extends LPA with the Community Overlap PRopagation Algorithm (COPRA), which removes nested communities and splits disconnected communities using a method similar to Raghavan et al. (Raghavan et al., 2007). Hafez et al. (Hafez et al., 2014) introduce a community detection algorithm using Bayesian Network and Expectation Minimization (BNEM), which includes a step to detect disconnected components within communities. If detected, disconnected components are assigned new community labels. This scenario arises when the network has more communities than the specified number $k$ in the algorithm. Hesamipour et al. (Hesamipour et al., 2022) propose a genetic algorithm for community detection, integrating similarity-based and modularity-based approaches. They use an MST-based representation to address issues like disconnected communities and ineffective mutations.
@@ 4.2. Explanation of BFS technique
We also attempted another parallel BFS-based algorithms to split disconnected communities. It begins by initializing the visited flags and BFS frontiers, using OpenMP to parallelize across vertices. Then, for each community, it selects a starting vertex and marks it as visited. Using a BFS approach, it iteratively explores neighbors within each community, updating each unvisited vertex’s community label, and marking it as visited if it shares the same community. The algorithm continues this process in a loop, advancing the BFS frontier until no new vertices are marked in a given iteration, at which point a reinitialization is triggered for untouched vertices in each community. This procedure is repeated until all vertices are processed. However, this algorithm offered lower performance than our BFS-based algorithm discussed above.
@@ 5.2. Comparing Performance of GSL-LPA
GSL-LPA demonstrates average speedups of $\times5$, 10,300×10,300$\times$10 , 300 ×, and 5.8×5.8$\times$5.8 × over FLPA, igraph LPA, and NetworKit LPA respectively. Specifically, on the sk-2005 graph, GSL-LPA identifies communities within $4.5$ seconds, achieving a processing rate of $844$ million edges/s. Furthermore, Figure 4(c) presents the modularity of communities obtained by each implementation. On average, GSL-LPA achieves modularity that is $7.1\%$ and $0.7\%$ higher than FLPA (particularly on road networks and protein k-mer graphs) and igraph LPA respectively. Additionally, it records a $3.6\%$ lower modularity compared to NetworKit LPA (especially on protein k-mer graphs). Finally, Figure 4(d) illustrates the fraction of disconnected communities obtained by each implementation. The absence of bars indicates the absence of disconnected communities. Communities identified by the GSL-LPA exhibit no disconnected communities. However, on average, FLPA, igraph LPA, and NetworKit LPA exhibit fractions of disconnected communities amounting to $2.3\times 10^{-2}$, $1.2\times 10^{-3}$, and $1.8\times 10^{-2}$, respectively, particularly noticeable on web graphs and social networks. Figure 7 depicts the comparison of GVE-LPA and GSL-LPA. This comparison is explained in detail in Section A.2.